MICROWAVE SPECTRA AND STRUCTURES OF OC-AuI AND ^{15}N_{2}-CuF

\underline{NICHOLAS R. WALKER}, SIMON G. FRANCIS, JOSHUA J. ROWLANDS and ANTHONY C. LEGON, School of Chemistry, University of Bristol, Cantock’s Close, Bristol, BS8 1TS, U.K..

A Fourier Transform microwave spectrometer coupled to a laser ablation source has been used to obtain the pure rotational spectra of OCAuI and ^{15}N_{2}-CuF. Both molecules are generated via the ablation of a metal rod in the presence of a halogen precursor and are stabilized by supersonic expansion. Both spectra are consistent with linear structures. The rotational constants (B_{0}, D_{J}) of OC-AuI have been measured to a high degree of precision permitting a determination of the molecular structure. The nuclear quadrupole coupling constants of the gold and iodine atoms are consistent with trends identified during studies of other OCMX species (M=Cu,Ag,Au; X=F,Cl,Br,I). An interim assignment of the spectrum of ^{15}N_{2}-CuF is made on the basis of lines from a single isotopomer. The nuclear quadrupole coupling constant of the copper atom in ^{15}N_{2}-CuF is of comparable magnitude to the value obtained in studies of OC-^{63}CuF. Lines in the pure rotational spectrum of ^{14}N_{2}-CuF are also presented.